An 800N elevator hangs by a steel cable for which allowable stress is 1.2 ×10^8 n/m².what is the minimum diameter required if the elevator accelerates upwards at 1.5 m/s²


1
Expert's answer
2021-03-04T07:34:14-0500

Let T be the tension in the cable.


Resolving upwards on the elevator gives: "T - 800 = 80 \times 1.5" as (F = ma)


Solve for T: "T = 120 + 800"


"T = 920 N"

Divide this by the stress to get the cross-sectional area required:


"A = \dfrac{920}{1.2\times 10^8}"

= "766.66\times 10^{-8}" "m^2"



Divide by pi to get "r^2"


"r^2 = 244.161\times 10^{-8}"


"r = 15.625\times 10^{-4} m"


Multiply by 2 to get the diameter


Hence, "d = 31.25\times 10^{-4}" m